Solve the following system by substitution.
-x+3y=6
y=1/3x+2

Answer by mananth(16946)

-1 x + 3 y = 6 .............1
1 y = 2 + 0.33 x ---------2
/ 1 =
y = 2 + 0.33 x
Substitute y in (1)
-1 x + 3 ( 2 + 0.33 x ) = 6
-1 x + 6 + 1 x = 6
-1 x + 1 x = 6 -6
0 x = 0
the lines are coinciding with each other. there are innumerable solutions.